OpenStudy (anonymous):

How many atoms are in 2.70 moles of iron (Fe) atoms? Select one: a. 4.48 × 10-24 atoms Fe b. 1.63 × 1024 atoms Fe c. 2.23 × 1023 atoms Fe d. 6.02 × 1023 atoms Fe

OpenStudy (anonymous):

@ash2326 @Nurali

OpenStudy (anonymous):

B

OpenStudy (anonymous):

I was thinking that to but I just wanted to make sure it was correct

OpenStudy (anonymous):

Yep, you just multiply it by avogrados number

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

P = nL P = 2.70 * 6.02*10^23 P = 1.63 * 10^24 atoms